There were 34,000 people at a ball game in Los Angeles. The day's receipts were $226,000. How many people paid $13 for reserved seats and how many paid $5 for general admission?
Answer by checkley77(12569)

13x+5(34,000-x)=226,000
13x+170,000-5x=226,000
8x=226,000-170,000
8x=56,000
x=56,000/8
x=7,000 $13 tickets were sold.
34,000-7,000=27,000 $5 tickets sold.
Proof:
13*7,000+5*27,000=226,000
91,000+135,000=226,000
226,000=226,000